OpenStudy (anonymous):

if y(ln3)=2 and y=(e^2x) - 1/(e^2x) + C, what is C?

OpenStudy (anonymous):

okay let y=2 and x=ln3, sub these values in to get C

OpenStudy (anonymous):

I know ho to do that but what is e^2ln3 ?

OpenStudy (anonymous):

e^(2ln3)=e^(ln3^2)=e^(ln9)=9

OpenStudy (anonymous):

ok so does e^lnx always equal x?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

thank you very much!